OpenStudy (anonymous):

Three functions are given below: f(x), g(x), and h(x). Explain how to find the axis of symmetry for each function, and rank the functions based on their axis of symmetry (from smallest to largest).

OpenStudy (anonymous):

f(x)=-4(x-8)^2+3 g(x) = 3x2 + 12x + 15

OpenStudy (anonymous):

@issaiahphillips

OpenStudy (anonymous):

The axis of symmetry is a vertical line with the equation of x = -b/2a. Start by getting f(x) into a simpler form. Something like ax^2+bx+c

OpenStudy (anonymous):

from my thoughts the biggest is f and the smallest is g and h is in the middle

OpenStudy (anonymous):

Yeah i got that its just finding the axis of symmetry that is getting me stuck. @issaiahphillips

OpenStudy (anonymous):

ooooh ok gotcha workin in it

OpenStudy (anonymous):

that's wrong @issaiahphillips I repeat, the axis of symmetry is a vertical line of form -b/2a so taking g as an example, the axis of symmetry is x=-12/6=-2

OpenStudy (aum):

The axis of symmetry is a vertical line that passes through the vertex. f(x) = -4(x - 8)^2 + 3 is already given in the vertex form. Vertex is (8, 3) So the axis of symmetry for f(x) is: x = 8

OpenStudy (aum):

g(x) is given in the standard form: ax^2 + bx + c The axis of symmetry is: x = -b/(2a)

OpenStudy (anonymous):

I just need h(x) i'm stuck on it @aum

OpenStudy (aum):

Identify the vertex from the graph.

OpenStudy (anonymous):

is the vertex (3,2)?

OpenStudy (aum):

Yes. The axis of symmetry is a vertical line that passes through the vertex.

OpenStudy (anonymous):

So then the axis of symmytry would be 3?

OpenStudy (aum):

The equation is: x = 3

OpenStudy (aum):

x = constant is the equation of a vertical line. y = constant is the equation of a horizontal line.

OpenStudy (anonymous):

Thank you! @aum

OpenStudy (aum):

You are welcome.